A system of equations is shown.
5x+y=9
10x−7y=−18

What are the values of x and y
Thanks so much for your help!!

Answer by macston(5194)

.
5x+y=9
y=9-5x Use this to substitute for y.
.
10x-7y=-18 Substitute for y from above.
10x-7(9-5x)=-18
10x-63+35x=-18
45x=45
x=1
.
y=9-5x=9-5=4
.
ANSWER: x=1 and y=4
.
CHECK:
5x+y=9
5(1)+4=9
9=9
.
10x-7y=-18
10(1)-7(4)=-18
10-28=-18
-18=-18